# A Tverberg-type problem of Kalai: Two negative answers to questions of Alon and Smorodinsky, and the power of disjointness

### v1

@@ Abstract
  • We show that $F_{2}(2,s,s)=O(s\log s)$ by performing controlled planar geometric transformations, and constructing an auxiliary graph whose planarity yields the upper bound.
@@ 1.3 Our contributions
Our results settle both questions of Alon and Smorodinsky in the negative. In the planar case, we present an elementary yet previously unnoticed construction that already forces quadratic growth.
@@
For large $s$, this lower bound matches the general upper bound of Alon and Smorodinsky up to a multiplicative factor of $\log s$. A refined version actually yields a slightly stronger bound $f_{2}(2,s,s)>2s^{2}$, and correspondingly $f_{r}(d,s,\ldots,s)>2s^{r}$ for $d\geq 2r-2$. Since the proof of $f_{2}(2,s,s)>2s^{2}$ might become considerably longer and obscure the elegance of the core idea, to preserve the simplicity of the argument, we will present this refined construction in Section 6.
@@
Regarding Question 1.2(2), Alon and Smorodinsky [2] showed that for fixed $r$, $f_{r}(d,s,\ldots,s)\leq\textup{Poly}(s,d)$; by (1) the polynomial bound also holds for $F_{r}(d,s,\ldots,s)$. In the complementary regime, according to the proof of Theorem 1.7, one can see that for fixed $d$, $F_{r}(d,s,\ldots,s)$ is also bounded by $O_{d}(r^{3}\log{r}\cdot s^{2d+3})$. Moreover, when both $r$ and $d$ are fixed, Theorem 1.7 improves the bound $O_{r,d}(s^{r})$ from Theorem 1.1 to $O_{r,d}\big(s^{(1-\varepsilon_{d})r+1}\big)$ for $\varepsilon_{d}=\frac{1}{2^{d}(d+1)}$ and $O_{r,d}\big(s^{2d+3}\big)$. It is not hard to see the bound $O_{r,d}(s^{2d+3})$ performs much better when $r$ is larger than $3d$.
@@ 2.5 Disjointness
$\angle(\bm{u}_{k}-\bm{y}_{k},\bm{u}_{k}-\bm{u}_{\ell})=\frac{\pi}{2}+\frac{m\pi}{s}\in[\frac{s+2}{2s}\pi,\pi],$
@@
$\angle(\bm{u}_{k},\bm{u_{\ell}})=\frac{2m\pi}{s}\in[\frac{2\pi}{s},\pi].$
@@
$\|\bm{u}_{k}-\bm{u}_{\ell}\|=2\sin{\frac{\angle(\bm{u}_{k},\bm{u_{\ell}})}{2}}\in[2\sin{\frac{\pi}{s}},2].$ (8)
@@ 4.1 Tools and auxiliary results
A graph is planar if it admits an embedding in the plane in which edges meet only at common endpoints. The following result about planar graphs is folklore.
Theorem 4.4.
Let $G$ be a simple planar graph with $n$ vertices and $m$ edges. If $n\geq 3$, then $m\leq 3n-6.$
@@
Our first ingredient is building the following relation between $F_{2}(2,s,s)$ and the above function.
@@
We the derive the following sharp linear estimate for $g(a)$. Then combining Theorem 4.8 with the reduction above immediately yields Theorem 1.6.
Theorem 4.8.
For every integer $a\geq 3$, $g(a)\leq 6a-12$.
Proof of Theorem 4.8.
To show $g(a)\leq g$, it suffices to show that for given pairwise disjoint convex sets $D_{1},\ldots,D_{a}$ in $\mathbb{R}^{2}$, there exists a separating system $P$ for $\mathcal{D}=\{D_{1},\ldots,D_{a}\}$ with $\textup{Fac}(P)\leq g.$ We begin with a brief preparation on separating systems.
Incidences in separating systems.
Let $P:=\{P_{1},\ldots,P_{m}\}$ be a separating system. We call a nonempty set $A\subset\mathbb{R}^{2}$ an incidence if there exist distinct $i,j\in[m]$ such that $A=\partial{P_{i}}\cap\partial{P_{j}}$ and $A$ is contained in the relative interior of some facet of $P_{i}$ or $P_{j}$. Therefore, we have two types of incidences:
  • Type 1. $A$ forms a non-degenerate closed line segment, or a closed ray, or a line.
  • Type 2. $A=\{\bm{a}\}$ is a singleton.
In particular, if $\partial{P_{i}}\cap\partial{P_{j}}=\{\bm{b}\}$ and $\bm{b}$ is a vertex of both $P_{i}$ and $P_{j}$, then we do not regard it as an incidence of the separating system $P$. See Fig. 4.1 for an illustration.
$\sqrt{}$
$\sqrt{}$
$\times$
Figure 4.1: The first two figures illustrate the two types of incidences in our definition, whereas the configuration in the third figure does not constitute an incidence.
Now, formally, we denote by $\mathcal{I}(P)$ the set of all incidences of $P$, moreover, we denote by $\mathcal{I}_{1}(P)$ and $\mathcal{I}_{2}(P)$ the sets of all incidences of $P$ of Type 1 and Type 2, respectively. Since for each unordered pair $(i,j)\in\binom{[m]}{2}$, $\partial P_{i}\cap\partial P_{j}=P_{i}\cap P_{j}$ is convex, each unordered pair contributes at most one incidence, which yields that $|\mathcal{I}(P)|\leq\binom{|P|}{2}$.
Selection of the desired separating system.
Having set up these definitions, we now select an extremal separating system that will drive the argument. Recall that $\mathcal{P}_{\mathcal{D}}$ is the nonempty collection of all separating systems for $\mathcal{D}=\{D_{1},\ldots,D_{a}\}$. Let $p_{\textup{max}}=\max\{|\mathcal{I}(P)|:P\in\mathcal{P}_{\mathcal{D}}\}$ and $p^{\prime}_{\textup{max}}:=\max\{|\mathcal{I}_{1}(P)|:P\in\mathcal{P}_{\mathcal{D}},\penalty 10000\ |\mathcal{I}(P)|=p_{\textup{max}}\}$, and we choose $Q=\{Q_{1},\ldots,Q_{a}\}$ such that
$|\mathcal{I}(Q)|=p_{\textup{max}},\qquad|\mathcal{I}_{1}(Q)|=p^{\prime}_{\textup{max}},$ (13)
and
$\mathrm{Fac}(Q)=\min\bigl\{\mathrm{Fac}(P):P\in\mathcal{P}_{\mathcal{D}},\ |\mathcal{I}(P)|=p_{\textup{max}}\ \textup{and}\ |\mathcal{I}_{1}(P)|=p^{\prime}_{\textup{max}}\bigr\}.$ (14)
Properties of selected separating system $Q$
Fix the above choice of $Q=\{Q_{1},\ldots,Q_{a}\}$ such that both (13) and (14) hold, and we then show the following key property. For a facet $e$ of some convex polyhedron, the open facet is the relative interior of $e$.
Claim 4.9.
For each $i\in[a]$ and each open facet $e$ of the convex polyhedron $Q_{i}$, there exists $j\neq i$ such that $e\cap\partial Q_{j}\neq\emptyset$. Equivalently, the relative interior of every facet of $\partial Q_{i}$ contains an incidence of $\mathcal{I}(\mathcal{Q})$.
Proof of claim.
It suffices to consider the case where $e$ is a line segment, since the same argument applies to the cases where $e$ is a ray or a line.
Fix $i\in[a]$ and a facet $\bm{a}\bm{b}\subseteq\partial Q_{i}$ with open facet $e=(\bm{a},\bm{b})$, and let $L_{e}$ be the line containing $e$. Let $W$ be the open exterior wedge at $e$, namely the open region bounded by $e$ and the two rays obtained by extending the adjacent facets of $\partial Q_{i}$ outward at $\bm{a}$ and $\bm{b}$, see Fig. 4.2 for the two possible shapes of $W$.
[…]
Figure 4.2: There are two possible shapes of $W$.
Assume, toward a contradiction, that $e\cap\partial Q_{j}=\emptyset$ for every $j\neq i$. We divide our argument into two cases, depending on whether $W$ intersects another polyhedron.
Case 1.
If $W\cap\bigl(\bigcup_{j\neq i}Q_{j}\bigr)\neq\emptyset,$ then we define
$d\;:=\;\inf\Bigl\{\mathrm{dist}(\bm{x},L_{e}):\ \bm{x}\in W\cap\bigg(\bigcup_{j\neq i}\partial Q_{j}\bigg)\Bigr\}\in[0,\infty).$
If $d>0$, then translate $L_{e}$ outward parallel to itself within $W$ until the translated segment $\overline{L^{\prime}\cap W}$ first meets $\partial Q_{k}$ for some $k$ such that $Q_{k}\cap W\neq\emptyset$. Let the two endpoints of $L^{\prime}\cap W$ be $\bm{a}^{\prime}$ and $\bm{b}^{\prime}$. By interior-disjointness and the definition of $d$, there are three situations:
  1. (1)
    $L^{\prime}$ meets $\partial{Q_{k}}$ at some point $\bm{x}\in{W}$, see Fig. 4.5;
  2. (2)
    $L^{\prime}\cap\partial{Q_{k}}$ is a segment, see Fig. 4.5;
  3. (3)
    $L^{\prime}\cap\partial{Q_{k}}$ is exactly $\bm{a}^{\prime}$ or $\bm{b}^{\prime}$, see Fig. 4.5.
$L_{e}$$Q_{k}$$\bm{a}$$\bm{b}$
$\rightarrow$
$L^{\prime}$$\bm{x}$$Q_{k}$
Figure 4.3: Case 1 (1), $d>0$, $L^{\prime}$ meets $\partial{Q_{k}}$ at some point $\bm{x}\in{W}$.
$L_{e}$$Q_{k}$$\bm{a}$$\bm{b}$
$\rightarrow$
$L^{\prime}$$Q_{k}$
Figure 4.4: Case 1 (2), $d>0$, $L^{\prime}\cap\partial{Q_{k}}$ is a segment.
$L_{e}$$\bm{a}$$\bm{b}$
$\rightarrow$
$\bm{a}^{\prime}$$\bm{b}^{\prime}$$L^{\prime}$
Figure 4.5: Case 1 (3), $d>0$, $L^{\prime}\cap\partial{Q_{k}}$ is exactly $\bm{a}^{\prime}$ or $\bm{b}^{\prime}$.
When the first or the second situations occurs, replacing the facet $\bm{a}\bm{b}$ of $Q_{i}$ by $L^{\prime}\cap\overline{W}$ yields a new separating system $Q^{\prime}$ with $|\mathcal{I}(Q^{\prime})|>|\mathcal{I}(Q)|$, contradicting the maximality of $|\mathcal{I}(Q)|$. When the third case occurs, that means, $L^{\prime}\cap\partial{Q_{k}}$ is exactly $\bm{a}^{\prime}$ or $\bm{b}^{\prime}$, and we can treat it as the case of $d=0$.
Next, it suffices to consider the case of $d=0$. In this case, some $Q_{k}$ meets the closure $\overline{e}$ at $\bm{a}$ or $\bm{b}$ by the assumption that $e\cap Q_{k}=\emptyset$ for each $k$. Let
$\mathcal{F}_{\bm{b}}:=\{f:f\ \textup{is\ a\ facet\ of}\ Q_{j}\ \textup{for\ some\ }j\in[a],f\cap e=\{\bm{b}\}\ \textup{and\ }f\cap W\neq\emptyset\}$
and
$\mathcal{F}_{\bm{a}}:=\{f:f\ \textup{is\ a\ facet\ of}\ Q_{j}\ \textup{for\ some\ }j\in[a],f\cap e=\{\bm{a}\}\ \textup{and\ }f\cap W\neq\emptyset\}.$
Without loss of generality, we can assume that both of $\mathcal{F}_{\bm{a}}$ and $\mathcal{F}_{\bm{b}}$ are nonempty. For each $f\in\mathcal{F}_{\bm{a}}\cup\mathcal{F}_{\bm{b}}$, we can arbitrarily select a point $\bm{w}(f)\in f\cap W$. Then let $f_{\bm{a}}\in\mathcal{F}_{\bm{a}}$ satisfy that the angle $\angle(\bm{a}-\bm{b},\bm{a}-\bm{w}(f_{\bm{a}}))$ is the smallest among all $f\in\mathcal{F}_{\bm{a}}$, and let $f_{\bm{b}}\in\mathcal{F}_{\bm{b}}$ satisfy that the angle $\angle(\bm{b}-\bm{a},\bm{b}-\bm{w}(f_{\bm{b}}))$ is the smallest among all $f\in\mathcal{F}_{\bm{b}}$. Then we denote by $T$ the region bounded by the ray supporting $f_{\bm{a}}$ with the endpoint $\bm{a}$, the ray supporting $f_{\bm{b}}$ with the endpoint $\bm{b}$, and $e$, see Fig. 4.6.
[…]
$\rightarrow$
$L_{\epsilon}$$T_{\epsilon}$
Figure 4.6: Case 1, $d=0$.
Then for any $k\neq i$, if $\textup{dist}(Q_{k}\cap W,e)=0$, then $Q_{k}\cap\{\bm{a},\bm{b}\}\neq\emptyset$. In this case, without loss of generality, suppose that $\bm{a}\in Q_{k}$, noting that $\bm{a}\in\partial{Q_{k}}$, then by definition of $f_{\bm{a}}$, $Q_{k}$ is disjoint from the open cone area bounded by $f_{\bm{a}}$ and $e$, which also implies that $Q_{k}$ is disjoint from the region $T$. Therefore, for those $Q_{k}$ with $Q_{k}\cap T\neq\emptyset$, we can always assume $\textup{dist}(Q_{k}\cap W,L_{e})>0$.
We can choose a positive number $\varepsilon<\min\limits_{k:Q_{k}\cap T\neq\emptyset}\textup{dist}(Q_{k}\cap W,e)$, and we define
$T_{\varepsilon}:=\{\bm{x}\in T:\textup{dist}(\bm{x},L_{e})\leq\varepsilon\}.$
Since $\varepsilon$ is small enough, we can see $T_{\varepsilon}\cap Q_{k}=\emptyset$ for all $k\neq i$. Observe that $T_{\varepsilon}$ consists of the points that lie in the area bounded by $L_{e}$, $f_{\bm{a}}$, $f_{\bm{b}}$ and the line $L_{\varepsilon}$ parallel to $L_{e}$ with distance $\varepsilon$ such that $L_{\varepsilon}\cap Q_{i}=\emptyset$. Notice that $Q_{i}^{\prime}:=Q_{i}\cup T_{\varepsilon}$ forms a convex polyhedron. Let $Q_{j}$ be the polyhedron containing $f_{\bm{a}}$ as a facet. Then by replacing $Q_{i}$ with $Q_{i}^{\prime}$ in the original separating system, we obtain at least one extra incidence of Type 1, namely $Q_{i}^{\prime}\cap Q_{j}$, in the meanwhile the total number of incidences of new separating system does not decrease, which is a contradiction to (13).
Case 2.
If $W\cap\bigl(\bigcup_{j\neq i}Q_{j}\bigr)=\emptyset$, then we replace $Q_{i}$ with $Q_{i}\cup W$ in the original separating system (see Fig. 4.7). Then the number of facets decreases, which contradicts (14).
[…]
$\rightarrow$
$Q_{i}$
Figure 4.7: Case 2.
Therefore, in either case, we reach a contradiction. This finishes the proof. ∎
The construction of planar auxiliary graph.
Let $\mathcal{I}(Q)=\{I_{1},I_{2},\ldots,I_{q}\}$. We then construct an auxiliary graph to bound the number of incidences $\mathcal{I}(Q)$ from above. Let $G=(V,E)$ where $V=[a]$. For each $j\in[q]$, by definition there are at least one pair $(k,\ell)\in\binom{[a]}{2}$ such that $I_{j}=Q_{k}\cap Q_{\ell}$ and $I_{j}$ is not a vertex of $Q_{k}$. For each incidence $I_{j}$, we select exactly one such pair $(k_{j},\ell_{j})$ and form an edge $k_{j}\ell_{j}$ in $G$. By definition, we have $|E(G)|=|\mathcal{I}(Q)|.$
Claim 4.10.
$G$ is a simple planar graph.
Proof of claim.
By definition, $G$ is a simple graph. We next show that $G$ is planar. For each $k\in[a]$, select a point $\bm{x}_{k}\in Q_{k}^{\circ}.$ Note that each edge of $G$ corresponds to an incidence $I=Q_{i}\cap Q_{j}$. Moreover, we select a point $\bm{y}_{i,j}\in I$ by the following rules: if $I$ itself is a point, then we choose this point as $\bm{y}_{i,j}$, or otherwise, we can arbitrarily select one point in the relative interior of $I$. By this rule, it is easy to check that for each incidence $I=Q_{i}\cap Q_{j}$, the point $\bm{y}_{i,j}$ is unique.
We then draw piecewise-linear path $\bm{x}_{i}\bm{y}_{i,j}\bm{x}_{j}$ consisting of the segments $\bm{x}_{i}\bm{y}_{i,j}$} and $\bm{y}_{i,j}\bm{x}_{j}$}. Suppose that there exist $i_{1},j_{1},i_{2},j_{2}$ such that the piecewise-linear paths $\bm{x}_{i_{1}}\bm{y}_{i_{1},j_{1}}\bm{x}_{j_{1}}$ and $\bm{x}_{i_{2}}\bm{y}_{i_{2},j_{2}}\bm{x}_{j_{2}}$ intersect in the interior. Without loss of generality, assume that the line segments $\bm{x}_{i_{1}}\bm{y}_{i_{1},j_{1}}$ and $\bm{x}_{i_{2}}\bm{y}_{i_{2},j_{2}}$ intersect at a point other than $\bm{x}_{i_{1}}$, noting that $\bm{x}_{i_{1}}\bm{y}_{i_{1},j_{1}}\subseteq Q_{i_{1}}$ and $\bm{x}_{i_{1}}\bm{y}_{i_{1},j_{1}}\cap\partial{Q_{i_{1}}}=\{\bm{y}_{i_{1},j_{1}}\}$, similarly, $\bm{x}_{i_{2}}\bm{y}_{i_{2},j_{2}}\subseteq Q_{i_{2}}$ and $\bm{x}_{i_{2}}\bm{y}_{i_{2},j_{2}}\cap\partial{Q_{i_{2}}}=\{\bm{y}_{i_{2},j_{2}}\}$. If $i_{1}\neq i_{2}$, then the only possibility is that $\bm{y}_{i_{1},j_{1}}=\bm{y}_{i_{2},j_{2}}$, which implies that they correspond to the same edge of the graph. Otherwise $i_{1}=i_{2}$, by definition $\bm{x}_{i_{1}}\bm{y}_{i_{1},j_{1}}$ and $\bm{x}_{i_{2}}\bm{y}_{i_{2},j_{2}}$ intersect at a point other than $\bm{x}_{i_{1}}$, which implies that they are the same line segment and correspond to the same edge of $G$. This finishes the proof. ∎
$\bm{x}_{1}$$\bm{x}_{2}$$\bm{x}_{3}$$\bm{x}_{4}$$\bm{x}_{5}$$Q_{1}$$Q_{2}$$Q_{3}$$Q_{4}$$Q_{5}$$\bm{y}_{1,2}$$\bm{y}_{1,4}$$\bm{y}_{1,5}$$\bm{y}_{2,3}$$\bm{y}_{3,4}$$\bm{y}_{4,5}$
Figure 4.8: An illustration of planar drawing of the auxiliary graph $G$ with respect to $\{Q_{i}\}_{i\in[5]}$.
Completing the proof
By Claim 4.10 and Theorem 4.4, $|\mathcal{I}(Q)|=|E(G)|\leq 3a-6.$ Notice that for $i\in[a]$, each open facet of $Q_{i}$ contains at least one incidence and each incidence in $\mathcal{I}(Q)$ belongs to at most two such open facets, therefore, $|\mathrm{Fac}(Q)|\leq 2|\mathcal{I}(Q)|\leq 6a-12.$ By the definition of function $g(a)$, we have $g(a)\leq 6a-12$. This finishes the proof. ∎
@@ 6 Concluding remarks
In this paper we established two negative answers to the questions of Alon and Smorodinsky concerning Tverberg-type intersections of unions of convex sets. We proved that for all $r\geq 2$, $s\geq 1$, and $d\geq 2r-2$,
@@
which matches their general upper bound up to a logarithmic factor. Our construction combines a planar “ring configuration” with a gluing scheme on a high-dimensional torus. We further introduced the disjoint-union variant $F_{r}(d,s_{1},\ldots,s_{r})$, revealing that the disjointness drastically alters the quantitative behaviour: in particular, $F_{2}(2,s,s)=O(s\log s)$, and for general parameters we obtained upper bounds via a novel connection to hypergraph Turán theory.
@@
Finally, we present an improved construction to show $f_{2}(2,s,s)>2s^{2}$ as follows for interested readers and show an example in Fig. 6.1. To show $f_{2}(2,s,s)\geq 2s^{2}+1,$ we construct a set $P$ with $2s^{2}$ points using a similar approach as in Section 2.3 and Section 2.4, but with the following modification: For each $k\in[s]$, we place $2s$ points $\bm{z}_{k,1},\dots,\bm{z}_{k,2s}\in I_{k}$ in clockwise order along $S_{k}$. The set $P$ is then defined as
$P:=\{\bm{z}_{i,j}:i\in[s],j\in[2s]\},\qquad|P|=2s^{2}.$
Now consider an arbitrary bipartition $P=P_{1}\sqcup P_{2}$. For each row index $i\in[s]$ define the row-convex container
$C_{i}:=\operatorname{Conv}\{\bm{z}_{i,k}\in P_{1}:k\in[2s]\},$
and for each column index $j\in[s]$ define the column-convex container
$D_{j}:=\operatorname{Conv}\{\bm{z}_{k,2j-1},\bm{z}_{k,2j}\in P_{2}:k\in[s]\}.$
Finally put
$C:=\bigcup_{i=1}^{s}C_{i},\qquad D:=\bigcup_{j=1}^{s}D_{j}.$
By the definitions, $C$ and $D$ are $s$-convex sets. Moreover, it is not hard to check that $P_{1}\subseteq C$ and $P_{2}\subseteq D$. The proof of $C\cap D=\emptyset$ is slightly complicated, but follows by a similar argument to that in the proof of Proposition 2.5, so we omit the full details here.
$z_{11}$$z_{12}$$z_{13}$$z_{14}$$z_{15}$$z_{16}$$z_{17}$$z_{18}$$z_{21}$$z_{22}$$z_{23}$$z_{24}$$z_{25}$$z_{26}$$z_{27}$$z_{28}$$z_{31}$$z_{32}$$z_{33}$$z_{34}$$z_{35}$$z_{36}$$z_{37}$$z_{38}$$z_{41}$$z_{42}$$z_{43}$$z_{44}$$z_{45}$$z_{46}$$z_{47}$$z_{48}$
Figure 6.1: $f_{2}(2,4,4)>32$

### v2

@@ Abstract
  • We show that $F_{2}(2,s,s)=O(s\log s)$ by connecting it to a suitable line-separating function in the plane.
@@ 1.3 Our contributions
We settle both questions of Alon and Smorodinsky in the negative. In the planar case, we present an elementary yet previously unnoticed construction that already forces quadratic growth.
@@
For large $s$, this lower bound matches the general upper bound of Alon and Smorodinsky up to a multiplicative factor of $\log s$. A refined version actually yields a slightly stronger bound $f_{2}(d,s,s)>ds^{2}$, and correspondingly $f_{r}(d,s,\ldots,s)>(d-2r+4)\cdot s^{r}$ for $d\geq 2r-2$, which indicates the upper bound $O_{r}(ds^{r}\log{s})$ is very close to optimal when both of $s$ and $d$ are large enough. Since the proof of $f_{r}(d,s,\ldots,s)>(d-2r+4)\cdot s^{r}$ might become considerably longer and obscure the elegance of the core idea, to preserve the simplicity of the argument, we will present this refined construction in Section 6.
@@
Regarding Question 1.2(2), Alon and Smorodinsky [2] showed that for fixed $r$, $f_{r}(d,s,\ldots,s)\leq\textup{Poly}(s,d)$; by (1) the polynomial bound also holds for $F_{r}(d,s,\ldots,s)$. In the complementary regime, according to the proof of Theorem 1.7, one can see that for fixed $d$, $F_{r}(d,s,\ldots,s)$ is also bounded by $O_{d}(r^{3}\log{r}\cdot s^{2d+3})$. Moreover, when both $r$ and $d$ are fixed, Theorem 1.7 improves the bound $O_{r,d}(s^{r})$ of Theorem 1.1 to $O_{r,d}\big(s^{(1-\varepsilon_{d})r+1}\big)$ for $\varepsilon_{d}=\frac{1}{2^{d}(d+1)}$ and $O_{r,d}\big(s^{2d+3}\big)$. It is not hard to see the bound $O_{r,d}(s^{2d+3})$ performs much better when $r$ is larger than $3d$.
@@ 2.5 Disjointness
$\angle(\bm{u}_{k}-\bm{y}_{k},\bm{u}_{k}-\bm{u}_{\ell})=\frac{\pi}{2}+\frac{m\pi}{s}\in\bigg[\frac{s+2}{2s}\pi,\pi\bigg],$
@@
$\angle(\bm{u}_{k},\bm{u_{\ell}})=\frac{2m\pi}{s}\in\bigg[\frac{2\pi}{s},\pi\bigg].$
@@
$\|\bm{u}_{k}-\bm{u}_{\ell}\|=2\sin{\frac{\angle(\bm{u}_{k},\bm{u_{\ell}})}{2}}\in\bigg[2\sin{\frac{\pi}{s}},2\bigg].$ (9)
@@ 4.1 Tools and auxiliary results
It was shown in [14] that $g(a)$ grows linearly in $a$.
Theorem 4.6 (Theorem 2 [14]).
For every integer $a\geq 3$, $g(a)\leq 6a-9$.
We then establish the following relation between $F_{2}(2,s,s)$ and the above function.
@@
Then combining Theorem 4.6 with the reduction above immediately yields Theorem 1.6.
@@ 6 Concluding remarks
In this paper we establish two negative answers to the questions of Alon and Smorodinsky concerning Tverberg-type intersections of unions of convex sets. We prove that for all $r\geq 2$, $s\geq 1$, and $d\geq 2r-2$,
@@
which matches their general upper bound up to a logarithmic factor. Our construction combines a planar “ring configuration” with a gluing scheme on a high-dimensional torus. We further introduce the disjoint-union variant $F_{r}(d,s_{1},\ldots,s_{r})$, revealing that the disjointness drastically alters the quantitative behaviour: in particular, $F_{2}(2,s,s)=O(s\log s)$, and for general parameters we obtain upper bounds via a novel connection to hypergraph Turán theory.
As we have mentioned in Section 1.3, we can further improve the lower bound as
$f_{r}(d,s,\ldots,s)>(d-2r+4)s^{r},$
for any $s\geq 1$, $r\geq 2$ and $d\geq 2r-2$. This constitutes a further step toward closing the gap with the upper bound $O_{r}(ds^{r}\log{s})$ in Theorem 1.1 obtained by Alon and Smorodinsky when both of $s$ and $d$ are large. Here we describe the construction to show $f_{2}(d,s,s)>ds^{2}$ in details and the general lower bound follows from the same ideas as that in Section 3.
To achieve this, in $\mathbb{R}^{d}$, the original selection of $s^{2}$ points in $\mathbb{R}^{2}$ can be modified by replacing each circular arc with a spherical surface $\mathbb{S}^{d-1}$ in $\mathbb{R}^{d}$, and subsequently mapping each point on it to the $d$ vertices of a small (d1)(d-1)-dimensional regular simplex on this $\mathbb{S}^{d-1}$.
Formally, the construction of the $d\cdot s^{2}$ point set proceeds as follows. Following the approach used for selecting $s^{2}$ points in the proof of Theorem 1.3, we work in $\mathbb{R}^{d}$ and select $s$ spherical surfaces $\mathbb{S}^{d-1}$, denoted $\mathcal{S}_{k}\cong\mathbb{S}^{d-1}$ for $k\in[s]$, each centered at $\bm{y}^{\prime}_{k}=(\bm{y}_{k},0,\dots,0)\in\mathbb{R}^{d}$ with sufficiently large radius $R$ (R=M1R=M-1 satisfies a similar condition as that in (2)), where $\bm{y}_{k}\in\mathbb{R}^{2}$ is defined in (3). For each $i,j\in[s]$, we also denote $\bm{z}^{\prime}_{i,j}=(\bm{z}_{i,j},0,\dots,0)\in\mathbb{R}^{d}$, where $\bm{z}_{i,j}$ is defined in (6). Moreover, for each $i,j\in[s]$, we then define a hyperplane $\mathbb{H}_{i,j}$ intersecting $\mathcal{S}_{i}$ (not only at $\bm{z}^{\prime}_{i,j}$) such that the vector $\bm{z}^{\prime}_{i,j}-\bm{y}^{\prime}_{i}$ is orthogonal to $\mathbb{H}_{i,j}$, with the distance $\textup{dist}(\bm{z}^{\prime}_{i,j},\mathbb{H}_{i,j})=\delta$, where $\delta$ is chosen sufficiently small relative to $\|\bm{z}^{\prime}_{i,j}-\bm{z}^{\prime}_{i,j+1}\|$ for all $i,j\in[s]$.
For each $i,j\in[s]$, on the intersection $\mathbb{H}_{i,j}\cap\mathcal{S}_{i}$, we select $d$ points forming a (d1)(d-1)-dimensional regular simplex, which consists of the points $\bm{z}_{i,j,\ell}$ for $\ell\in[d]$. Then we define the set $P$ as
$P:=\{\bm{z}_{i,j,\ell}:i\in[s],\ j\in[s],\ \ell\in[d]\}.$
Note that $|P|=d\cdot s^{2}$. Now consider an arbitrary bipartition $P=P_{1}\sqcup P_{2}$. For each row index $i\in[s]$, define the row convex container
$C_{i}:=\operatorname{Conv}\{\bm{z}_{i,a,b}\in P_{1}:a\in[s],\ b\in[d]\},$
and for each column index $j\in[s]$, define the column convex container
$D_{j}:=\operatorname{Conv}\{\bm{z}_{a,j,b}\in P_{2}:a\in[s],\ b\in[d]\}.$
Finally, define the union sets $C:=\bigcup_{i=1}^{s}C_{i}$ and $D:=\bigcup_{j=1}^{s}D_{j}.$ By construction, $C$ and $D$ are $s$-convex sets. Furthermore, it is immediate from the definitions that $P_{1}\subseteq C$ and $P_{2}\subseteq D$. The proof that $C\cap D=\emptyset$ requires more complicated analysis but follows an argument similar to that in Proposition 2.5, here we give a brief sketch of the proof. For each $i,j\in[s]$, define that
$A_{i}:=\operatorname{Conv}\{\bm{z}_{i,a,b}:a\in[s],\ b\in[d]\},$
$B_{j}:=\operatorname{Conv}\{\bm{z}_{a,j,b}:a\in[s],\ b\in[d]\}.$
We have the following key observations:
  • The points $\{\bm{z}_{i,a,b}\}$ in $A_{i}$ and $\{\bm{z}_{a,j,b}\}$ in $B_{j}$ are in convex position, respectively.
  • $A_{i}\cap B_{j}=\operatorname{Conv}\{\bm{z}_{i,j,\ell}:\ell\in[d]\}$ since $A_{i}$ and $B_{j}$ lie in distinct closed half-spaces defined by $\mathbb{H}_{i,j}$.
Since the set $\{\bm{z}_{i,j,\ell}:\ell\in[d]\}$ constitutes the vertices of a (d1)(d-1)-dimensional regular simplex, for any partition $Z_{i,j}=Q_{1}\sqcup Q_{2}$ we have
$\operatorname{Conv}\{\bm{z}_{i,j,\ell}\in Q_{1},\ \ell\in[d]\}\cap\operatorname{Conv}\{\bm{z}_{i,j,\ell}\in Q_{2},\ \ell\in[d]\}=\emptyset.$
Consequently, $C_{i}\cap D_{j}=\emptyset$ for all $i,j\in[s]$.
We show an example for interested readers in Fig. 6.1.
$z_{1,1,1}$$z_{1,1,2}$$z_{1,2,1}$$z_{1,2,2}$$z_{1,3,1}$$z_{1,3,2}$$z_{1,4,1}$$z_{1,4,2}$$z_{2,1,1}$$z_{2,1,2}$$z_{2,2,1}$$z_{2,2,2}$$z_{2,3,1}$$z_{2,3,2}$$z_{2,4,1}$$z_{2,4,2}$$z_{3,1,1}$$z_{3,1,2}$$z_{3,2,1}$$z_{3,2,2}$$z_{3,3,1}$$z_{3,3,2}$$z_{3,4,1}$$z_{3,4,2}$$z_{4,1,1}$$z_{4,1,2}$$z_{4,2,1}$$z_{4,2,2}$$z_{4,3,1}$$z_{4,3,2}$$z_{4,4,1}$$z_{4,4,2}$
Figure 6.1: $f_{2}(2,4,4)>32$